OpenStudy (anonymous):

integrate with respect to x (e^(2x)-e^(-2x)) /(e^(2x) +e^(-2x))

OpenStudy (anonymous):

??????????/

OpenStudy (anonymous):

well now i got it ur answer is wrong

OpenStudy (anonymous):

i was thinking in any other way

OpenStudy (amistre64):

e^(2x)-e^(-2x) -------------- e^(2x) +e^(-2x) e^(4x)-1 --------- e^(4x) +1

OpenStudy (anonymous):

well there is a easy way than this one use subtitution XD directly XD

OpenStudy (amistre64):

there usually is :) i like to algebra them a few times to see if theres a simpler form hidden inside

OpenStudy (anonymous):

lol :P

OpenStudy (anonymous):

me too was doing in the same way :P wasted so many hours

OpenStudy (amistre64):

e^(4x)+1-2 ---------- e^(4x) +1 \[1-\frac2{(e^{2x})^2+1}\] looks tangible ...

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

u = e^2x; u' = 2 e^2x but then again lol

OpenStudy (anonymous):

hey i have one more question

OpenStudy (anonymous):

:P

OpenStudy (amistre64):

is it an easy one?

OpenStudy (anonymous):

well i think so

OpenStudy (anonymous):

integrate 1/[(sin(x-a)sin(x-b) ]

OpenStudy (anonymous):

wrt x

OpenStudy (amistre64):

i recall some trig stuff: cos(a+b) = cosa cosb - sina sinb -cos(a-b) = -cosa cosb - sina sinb ------------------------------ cos(a-b)/2 + cos(a+b)/2= sina sinb

OpenStudy (anonymous):

oh oh but in book they have done it only in one way

OpenStudy (amistre64):

let a = x-a, and b = x-b a+b = 2x-(a+b) a-b = b-a is a constant

OpenStudy (amistre64):

\[\frac{2}{cos(2x-k)+n}\]

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

with any luck, thats workable ....

OpenStudy (anonymous):

oh :) tahnx :)

OpenStudy (anonymous):

i wish if i could give u one more medal

OpenStudy (amistre64):

thnx, but these are just some ideas, i do not know if they produce something simpler to play with or not :)

OpenStudy (anonymous):

they will :P